OpenStudy (anonymous):

is the pair 3x+2y=6 and 3x+2y=9 parallel, perpendicular or neither? do I solve first?

OpenStudy (anonymous):

since the slope is same, so they are parallel

OpenStudy (anonymous):

slope meaning the equations?

OpenStudy (anonymous):

cause this graphing stuff I do not understand lol

OpenStudy (anonymous):

In each equation, solve for y. Set y on one side, everything else on the other side. y=mx+b m is the slope; the coefficient of x

OpenStudy (anonymous):

ok so what if the equation was 3x+2y=6 and 2y+3x=9 what would that be?

OpenStudy (anonymous):

Follow instructions above.

OpenStudy (anonymous):

ok thanks soo much. ive been trying to study for my exam ugh

OpenStudy (anonymous):

If the m is the same, for example, if m is 1/2 in each equation: they are parallel. If the m is what they call negative inverse, for example one is 1/2 the other is -2 they are perpendicular.